OpenStudy (anonymous):

Help, please? Which polynomial is a perfect square trinomial? 49x2 − 28x + 16 4a2 − 20a + 25 25b2 − 20b − 16 16x2 − 24x − 9

OpenStudy (welshfella):

find the square root of the first and last terms and try to combine them to get the middle term

OpenStudy (welshfella):

* first and last coefficients I should have said

OpenStudy (welshfella):

(a + b^2 = a^2 + 2ab + b^2 a and b have to such that they can form 2ab

OpenStudy (anonymous):

I want to say B but I'm not sure-

OpenStudy (welshfella):

the last 2 option are incorrect because there are no real square roots of -16 and -9

OpenStudy (welshfella):

what are the square roots of 2a^2 and 25

OpenStudy (welshfella):

* 4a^2 and 25?

OpenStudy (welshfella):

remember that 25 = 5*5 or -5*-5

OpenStudy (anonymous):

The square root of 25 is 5, but I'm not sure about 4a^2

OpenStudy (welshfella):

2a * 2a = 4a^2

OpenStudy (welshfella):

sqrt 25 is 5 OR -5

OpenStudy (anonymous):

Okay, that makes sense, so wait would the square root of 4a^2 be 2a?

OpenStudy (welshfella):

so if a = 2a and b = -5 what is the value of 2ab?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

So it would be 2a*5 since b=5?

OpenStudy (anonymous):

Or -5 I should say.

OpenStudy (welshfella):

no 2 * 2a * -5

OpenStudy (welshfella):

what does that come to?

OpenStudy (anonymous):

2a*25, right? Since you really can't do anything with the A right now because we don't know what that is-

OpenStudy (anonymous):

NONO NEVERMIND 2a*10

OpenStudy (welshfella):

2a * -10

OpenStudy (welshfella):

its NEGATIVE 5 here because the middle term is negative

OpenStudy (anonymous):

Oh, okay. That makes more sense.

OpenStudy (welshfella):

so what does 2*-10a simplify to?

OpenStudy (anonymous):

-20a?

OpenStudy (anonymous):

I think that's right at least-